Two different simple random samples are drawn from two different populations. The first sample consists of 40

people with 22 having a common attribute. The second sample consists of 2000

people with 1430 of them having the same common attribute. Compare the results from a hypothesis test of p1equals=p 2

​(with a 0.01 significance​ level) and a 99​% confidence interval estimate of p1−p 2

a) identify test statistic

b) identify critical value

c)What is the conclusion based on the hypothesis​ test?

d)the 99% confidence interval is
Answer #1

Answer

(A) p1 = x1/n1 = 22/40 = 0.55

p2 = x2/n2 = 1430/2000 = 0.715

p = (x1+x2)/(n1+n2)= (22+1430)/(40+2000) = 0.712

test statistic =

(B) z critical = -2.576 and 2.576 (using z table for 99% confidence level)

we have two critical values because we are testing for two tailed hypothesis

(C) z statistic is between z critical range, so we failed to reject the null hypothesis.

(D) Confidence interval =